# Understanding Linear Probing then Fine-tuning Language Models from NTK Perspective

### v1

@@ Abstract
The two-stage fine-tuning (FT) method, linear probing then fine-tuning (LP-FT), consistently outperforms linear probing (LP) and FT alone in terms of accuracy for both in-distribution (ID) and out-of-distribution (OOD) data. This success is largely attributed to the preservation of pre-trained features, achieved through a near-optimal linear head obtained during LP. However, despite the widespread use of large language models, the exploration of complex architectures such as Transformers remains limited. In this paper, we analyze the training dynamics of LP-FT for classification models on the basis of the neural tangent kernel (NTK) theory. Our analysis decomposes the NTK matrix into two components, highlighting the importance of the linear head norm alongside the prediction accuracy at the start of the FT stage. We also observe a significant increase in the linear head norm during LP, stemming from training with the cross-entropy (CE) loss, which effectively minimizes feature changes. Furthermore, we find that this increased norm can adversely affect model calibration, a challenge that can be addressed by temperature scaling. Additionally, we extend our analysis with the NTK to the low-rank adaptation (LoRA) method and validate its effectiveness. Our experiments with a Transformer-based model on natural language processing tasks across multiple benchmarks confirm our theoretical analysis and demonstrate the effectiveness of LP-FT in fine-tuning language models. Code is available at https://github.com/tom4649/lp-ft_ntk.
@@ 4.2 Training dynamics in the NTK regime
The proof of this proposition is included in the Appendix (Section 7.2.1). In our decomposition of the NTK matrix, the pre-train-effective component $\bm{P}(\bm{x},\bm{x}_{i})$ is a diagonal matrix and remains unchanged after LP, while the FT-effective component $\bm{F}(\bm{x},\bm{x}_{i})$ is not a diagonal matrix and does change after LP, resulting in distinct characteristics for these components. The Frobenius norm of the classifier weight matrix, $\|\bm{V}_{0}\|_{F}$, influences the balance between the pre-train-effective and FT-effective components because it affects only the FT-effective component. This indicates that the classifier weight norm $\|\bm{V}_{0}\|_{F}$ has a significant impact on the training dynamics of FT.
@@
This corollary shows that feature vectors for the samples in the orthogonal complement of training sample subspace are not updated. Therefore, given that pre-trained features have characteristics beneficial to downstream tasks, significant feature changes in FT, dependent on small training samples in LP, lead to poor generalization and OOD performance. The proof of this lemma can be found in the Appendix (Section 7.2.2).
@@ 4.4 Increase in the classifier weight norm
$\frac{\partial L(\bm{f})}{\partial\|\bm{v}_{k}\|}$ $=\sum_{i=1}^{N}\left(\sum_{y_{i}\neq k}[\bm{\sigma}_{\text{SM}}\left(\bm{f}(\bm{x_{i}})\right)]_{k}\|\bm{\phi}(\bm{x_{i}})\|\cos\tau_{ki}-\sum_{y_{i}=k}(1-[\bm{\sigma}_{\text{SM}}\left(\bm{f}(\bm{x_{i}})\right)]_{y_{i}})\|\bm{\phi}(\bm{x_{i}})\|\cos\tau_{y_{i}i}\right).$ (8)
@@ 4.5 Training process of LoRA
This proposition suggests that with high probability, the only difference of the NTK matrix between LoRA and standard FT is a scalar factor of the FT-effective component in the NTK matrix, and the scalar factor depends on the hyperparameters of LoRA. This implies that when the hyperparameters of LoRA are set appropriately, LoRA training is similar to standard FT training. This is consistent with the analysis by Malladi et al. [34], where the NTK matrix of LoRA and standard FT are close to each other. It is important to note that the proposition is also valid for LP-FT and LP-LoRA (LP then LoRA). The proof of this proposition is included in the Appendix (Section 7.2.3).
@@ 5 Numerical evaluation with transformer models
Details on the datasets, setup, and additional results, including performance evaluations for the experimental and practical application, are available in the Appendix (Sections 7.3 and 7.4).
@@ 5.2 Small feature changes during LP-FT and significant norm increase during LP
LP-FT achieves notable performance with Transformer-based language models, outperforming standard FT in both ID and OOD settings, as detailed in Appendix (Sections 7.4.1 and 7.4.3). To understand the underlying reasons for these results and validate small feature changes suggested by our analysis (Section 4.2), we analyzed changes in both the classifier and the features.

### v2

@@ Abstract
The two-stage fine-tuning (FT) method, linear probing (LP) then fine-tuning (LP-FT), outperforms linear probing and FT alone. This holds true for both in-distribution (ID) and out-of-distribution (OOD) data. One key reason for its success is the preservation of pre-trained features, achieved by obtaining a near-optimal linear head during LP. However, despite the widespread use of large language models, there has been limited exploration of more complex architectures such as Transformers. In this paper, we analyze the training dynamics of LP-FT for classification tasks on the basis of the neural tangent kernel (NTK) theory. Our analysis decomposes the NTK matrix into two components. This decomposition highlights the importance of the linear head norm alongside the prediction accuracy at the start of the FT stage. We also observe a significant increase in the linear head norm during LP, which stems from training with the cross-entropy (CE) loss. This increase in the linear head norm effectively reduces changes in learned features. Furthermore, we find that this increased norm can adversely affect model calibration, which can be corrected using temperature scaling. Additionally, we extend our analysis with the NTK to the low-rank adaptation (LoRA) method and validate its effectiveness. Our experiments using a Transformer-based model on multiple natural language processing datasets confirm our theoretical analysis. Our study demonstrates the effectiveness of LP-FT for fine-tuning language models. Code is available at https://github.com/tom4649/lp-ft_ntk.
@@ 4.2 Training dynamics in the NTK regime
The proof of this proposition is included in the Appendix (Section A.2.1). In our decomposition of the NTK matrix, the pre-train-effective component $\bm{P}(\bm{x},\bm{x}_{i})$ is a diagonal matrix and remains unchanged after LP, while the FT-effective component $\bm{F}(\bm{x},\bm{x}_{i})$ is not a diagonal matrix and does change after LP, resulting in distinct characteristics for these components. The Frobenius norm of the classifier weight matrix, $\|\bm{V}_{0}\|_{F}$, influences the balance between the pre-train-effective and FT-effective components because it affects only the FT-effective component. This indicates that the classifier weight norm $\|\bm{V}_{0}\|_{F}$ has a significant impact on the training dynamics of FT.
@@
This corollary shows that feature vectors for the samples in the orthogonal complement of training sample subspace are not updated. Therefore, given that pre-trained features have characteristics beneficial to downstream tasks, significant feature changes in FT, dependent on small training samples in LP, lead to poor generalization and OOD performance. The proof of this lemma can be found in the Appendix (Section A.2.2).
@@ 4.4 Increase in the classifier weight norm
$\frac{\partial L(\bm{f})}{\partial\|\bm{v}_{k}\|}$ $=\sum_{i=1}^{N}\left(\sum_{k\neq y_{i}}[\bm{\sigma}_{\text{SM}}\left(\bm{f}(\bm{x_{i}})\right)]_{k}\|\bm{\phi}(\bm{x_{i}})\|\cos\tau_{ki}-\sum_{k=y_{i}}(1-[\bm{\sigma}_{\text{SM}}\left(\bm{f}(\bm{x_{i}})\right)]_{y_{i}})\|\bm{\phi}(\bm{x_{i}})\|\cos\tau_{y_{i}i}\right).$ (8)
@@ 4.5 Training process of LoRA
This proposition suggests that with high probability, the only difference of the NTK matrix between LoRA and standard FT is a scalar factor of the FT-effective component in the NTK matrix, and the scalar factor depends on the hyperparameters of LoRA. This implies that when the hyperparameters of LoRA are set appropriately, LoRA training is similar to standard FT training. This is consistent with the analysis by Malladi et al. [34], where the NTK matrix of LoRA and standard FT are close to each other. It is important to note that the proposition is also valid for LP-FT and LP-LoRA (LP then LoRA). The proof of this proposition is included in the Appendix (Section A.2.3).
@@ 5 Numerical evaluation with transformer models
Details on the datasets, setup, and additional results, including performance evaluations for the experimental and practical application, are available in the Appendix (Sections A.3 and A.4).
@@ 5.2 Small feature changes during LP-FT and significant norm increase during LP
LP-FT achieves notable performance with Transformer-based language models, outperforming standard FT in both ID and OOD settings, as detailed in Appendix (Sections A.4.1 and A.4.3). To understand the underlying reasons for these results and validate small feature changes suggested by our analysis (Section 4.2), we analyzed changes in both the classifier and the features.